Question 1 Your answer is partially correct. Try again A 54.7-kg person, running horizontally with a velocity of +2.54 m/s, jumps onto a 18.5-kg sled that is initially at rest. (a) Ignoring the effects of friction during the collision, find the velocity of the sled and person as they move away. (b) The sled and person coast 30.0 m on level snow before coming to rest. What is the coefficient of kinetic friction between the sled and the snow? (a) Num umber Number L0.0242 1.92 : Units |Lm/s $ Units No units
Answer #1

Let H combined eloeit is V miu V: 51.7x54 54.7 +8.5 I. 89 8 = 1.9 m/s Lat Ketan datiam is a a = 0.06 m/s = 4.4 N 17.36 = 0.00